# Systematic Bayesian Optimization for Atomic Structure Calculations of Heavy Elements

### v1

@@ Abstract
This study presents a novel optimisation technique for atomic structure calculations using the Flexible Atomic Code, focussing on complex multielectron systems relevant to -process nucleosynthesis and kilonova modelling. We introduce a method to optimise the fictitious mean configuration used in the Flexible Atomic Code, significantly improving the accuracy of calculated energy levels and transition properties for lanthanide and actinide ions. Our approach employs a Sequential Model-Based Optimisation algorithm to refine the fictitious mean configuration, iteratively minimising the discrepancy between calculated and experimentally determined energy levels. We demonstrate the efficacy of this method through detailed analyses of Au II, Pt II, Pr II, Pr III, Er II, and Er III, representing a broad range of atomic configurations. The results show substantial improvements in the accuracy of the calculated energy levels, with average relative differences to the NIST data reduced from 20-60% to 10% or less for the ions studied. Transition wavelength calculations exhibit exceptional agreement with experimental data, with about 90% of the calculated values falling within 10% of measurements for Pr and Er ions. While improvements in transition probability calculations are observed, the calculated transition probabilities (log() values) still show significant discrepancies compared to the experimental data, with root mean square deviations of approximately 1.1-1.4 dex for Pr and Er ions. We extend our optimisation technique to systematic calculations of singly and doubly ionised lanthanides, achieving accuracies comparable to or surpassing those of ab-initio atomic structure codes. The method’s broad applicability across the lanthanide series demonstrates its potential for enhancing opacity calculations and spectral modelling in astrophysical contexts.
@@ III FMC optimisation procedure
To evaluate different potentials produced by , we employ a sequential model-based optimisation (SMBO) approach. At each iteration, generates a potential from a given set of . This potential is used in the FAC calculations to produce energy levels, which are then compared to the experimental data. Although various statistical measures were considered for this comparison, the limited and nonuniform availability of experimental uncertainties for lanthanide and actinide energy levels made such approaches impractical. Instead, we found that a weighted root mean square deviation (WRMSD) provided the most reliable convergence:
where is the difference between the calculated and reference energy levels. The Boltzmann factor accounts for the thermal population distribution under local thermodynamic equilibrium conditions, which makes optimisation particularly relevant for kilonova modelling. While other metrics such as weighted average deviation and level density were tested, WRMSD provided better convergence with fewer iterations. Using A-values or oscillator strengths as evaluation metrics was considered but rejected as it would require computing these values at each iteration, significantly increasing the computational time.
@@
This methodology is general and can be adapted to other atomic structure codes, with the flexibility to modify the loss function for different optimisation needs. The systematic evaluation of the predicted atomic data’ sensitivity to the optimised mean local potential provides valuable insights into the reliability and precision of the results.
@@ IV.1.3 Energy Levels
[…]
@@ IV.2.2 Computational procedure
The basis space for Pr ii and Pr iii was expanded to include all major configurations up to the ionisation energy. For Pr ii and Pr iii, it was found that considering SD excitations from the ground configuration ( for Pr ii and for Pr iii) up to , , , , effectively covered the most relevant configurations that contribute directly to opacity [35]. These configurations were used in the optimisation procedure in order to determine the optimal FMC for these ions. Hence, we label this space as Optimisation Space (OPS). To further enhance the convergence of the results, additional configurations were included based on SD excitations extending up to principal quantum number ) angular momentum . The space was increased in multiple steps from a base ground configuration (GC), with each space used labelled as a different CI space (CIS). The scheme for the full set of calculations was as follows.
@@ IV.2.3 Energy Levels
The average relative difference from the NIST data for all identified levels of Pr ii and Pr iii for the different models calculated with FAC is shown in Fig. 6. The “No Opt.” model corresponds to the OPS configuration model that does not use the FAC potential optimisation, whereas the other models incorporate this optimisation. The optimisation leads to a substantial reduction in the relative difference, far exceeding the improvements achieved by CI alone. Additionally, the convergence of results across different CI spaces (CIS1, CIS2, and CIS3) indicates that the optimisation process stabilises the calculations, ensuring reliable and consistent results. The results for the energy levels of the first 20 computed levels are shown in Tables 1 and 2.
@@ IV.2.4 Transitions
To evaluate the accuracy of our calculations for transition probabilities and wavelengths in Pr ii and Pr iii, we compared our results with recent experimental data from Ferrara et al. (2024) [72]. Their study provides valuable spectroscopic data for several high- elements, including praseodymium, covering the 3700-10000 Å range using high-resolution échelle spectroscopy. Figure 7 presents a comparison between the calculated values from our CIS3 model and the experimental data. The experimental values are derived using two methods: Boltzmann plots and cascade calculations. It’s important to note that while cascade method results are generally reliable within the limits of adopted literature values, the Boltzmann method is only valid for transitions in thermodynamic equilibrium. When available, we also evaluated our data against data from the Vienna Atomic Line Database (VALD) [73]. As evident from Fig. 7, there is considerable scatter in the values, particularly for weaker transitions (). The agreement between our calculations and the experimental data improves for stronger transitions, but discrepancies persist across the entire range. Quantitatively, we observe a root mean square deviation (RMSD) of approximately 1.1 dex between our calculated values and the experimental data. This level of disagreement is not uncommon in atomic structure calculations for complex ions like Pr ii and Pr iii, especially given the challenges in accurately representing the electron correlations in these systems. In contrast, Fig. 8 shows excellent agreement between our calculated wavelengths and the experimental values. Most of the calculated wavelengths fall within 10% of the experimental data, as indicated by the points between the inner dotted lines. Nearly all calculated wavelengths are within 20% of the experimental values, falling between the dashed lines. Quantitatively, we find that approximately 80% of our calculated wavelengths deviate less than 20% from the experimental values and 96% are within 10%. This disparity in accuracy between the transition probabilities and wavelengths is a common feature in atomic structure calculations. Wavelengths, which are primarily determined by energy level differences, are often more accurately predicted than transition probabilities, which depend on the precise details of the wavefunctions.
@@ IV.3.1 Available data
Building on this work, Goldschmidt (1963) [78] computed all 24 levels of the 6s configuration by fitting the Slatter parameters using the initial eight experimentally determined levels as a reference. Subsequently, Marquet & Davis (1965) [79] identified 10 levels of the ground configuration using wavelengths in the visible and near-ultraviolet regions. Their Zeeman structure measurements of 132 transitions enabled Sluis & McNally (1970) [80] to identify 12 of the lowest ground configuration levels, which exhibited slightly higher energies than those calculated by Goldschmidt (1963) with = 30. Additionally, Sluis & McNally (1970) reported 90 odd levels, but these were excluded from comparisons due to a lack of configuration identification and assignment.
@@ IV.3.4 Transitions
In contrast, Fig. 13 demonstrates excellent agreement between our calculated wavelengths and the experimental values for Er ii. Most of the calculated wavelengths fall well within 10% of the experimental data, as indicated by the points that lie between the inner dotted lines. Virtually all calculated wavelengths are within 20% of the experimental values, falling between the dashed lines. Quantitatively, we find that approximately 85% of our calculated wavelengths deviate by less than 10% from the experimental values, and over 95% are within 20%.
The stark contrast in accuracy between the transition probabilities and the wavelengths observed for Er ii mirrors our findings for Pr ii. This discrepancy can be attributed to the nature of our optimisation procedure, which is primarily designed to minimise differences in energy levels when compared to experimental data. While this approach typically leads to more accurate wavefunctions, these results clearly demonstrate its limitations. The high accuracy in wavelength predictions, which are primarily determined by energy level differences, is a direct consequence of this optimisation strategy. However, persistent discrepancies in oscillator strengths suggest that our current approach may not fully capture the intricacies of the wavefunctions required for accurate transition probability calculations. This indicates that including a larger basis set in our calculations may be necessary to improve the accuracy of oscillator strengths, as these are more sensitive to the fine details of the electronic wavefunctions.
@@ VI Conclusions
However, persistent challenges in calculating accurate transition probabilities underscore the need for further refinement of our approach. Future work should focus on expanding the basis set to potentially improve the accuracy of oscillator strengths and transition probabilities. Although computationally demanding, investigating the inclusion of core-core and core-valence correlations may further enhance the accuracy of energy level and transition calculations. The demonstrated success of our optimisation approach in improving energy level calculations is particularly valuable as it enables more reliable level identification. This enhanced identification capability provides a solid foundation for future work involving detailed calibration of theoretical energy levels against experimental data [89].

### v2

@@ Abstract
This study presents a novel optimisation technique for atomic structure calculations using the Flexible Atomic Code, focussing on complex multielectron systems relevant to -process nucleosynthesis and kilonova modelling. We introduce a method to optimise the fictitious mean configuration used in the Flexible Atomic Code, significantly improving the accuracy of calculated energy levels and transition properties for lanthanide and actinide ions. Our approach employs a Sequential Model-Based Optimisation algorithm to refine the fictitious mean configuration, iteratively minimising the discrepancy between calculated and experimentally determined energy levels. We demonstrate the efficacy of this method through detailed analyses of Au II, Pt II, Pr II, Pr III, Er II, and Er III, representing a broad range of atomic configurations. The results show substantial improvements in the accuracy of the calculated energy levels, with average relative differences to the NIST data reduced from 20-60% to 10% or less for the ions studied. Transition wavelength calculations exhibit good agreement with experimental data, with about 90% of the calculated values falling within 10% of measurements for Pr and Er ions. While improvements in transition probability calculations are observed, the calculated radiative transition probabilities (log() values) still show significant discrepancies compared to the experimental data, with root mean square deviations of approximately 1.1-1.4 dex for Pr and Er ions. We extend our optimisation technique to systematic calculations of singly and doubly ionised lanthanides, achieving accuracies comparable to or surpassing those of ab-initio atomic structure codes. The method’s broad applicability across the lanthanide series demonstrates its potential for enhancing opacity calculations and spectral modelling in astrophysical contexts.
@@ III FMC optimisation procedure
To evaluate different potentials produced by , we employ a sequential model-based optimisation (SMBO) approach. At each iteration, generates a potential from a given set of . This potential is used in the FAC calculations to produce energy levels, which are then compared to the experimental data. We explored various statistical measures for this comparison, including mean absolute deviation, weighted average deviation, and level density. However, the limited and nonuniform availability of experimental uncertainties for lanthanide and actinide energy levels made some approaches impractical. After testing, we found that a weighted root mean square deviation (WRMSD) provided the most reliable convergence
where is the difference between the calculated and reference energy levels. The Boltzmann factor with K accounts for the thermal population distribution under LTE conditions, giving higher weight to lower-lying levels that contribute most significantly to opacity in kilonova. We note that as long as the temperature is representative of an LTE regime in kilonovae, small variations in the specific value have minimal impact on the optimization results. Using A-values or oscillator strengths as evaluation metrics was considered but rejected as it would require computing these values at each iteration, significantly increasing the computational time.
@@
This optimization methodology is general and can be adapted to other atomic structure codes that rely on local central potentials or scaling parameters that require user input. This includes codes such as HULLAC and AUTOSTRUCTURE, where scaling parameters are used. While our specific implementation targets the fictitious mean configuration in FAC, the core principles, employing Bayesian optimization to minimize the discrepancy between calculated and reference energy levels, could be applied to any code where the potential or parameters can be systematically adjusted to improve agreement with reference data. The effectiveness of this approach would be greatest in codes where the calculated energy levels are sensitive to user-defined input parameters that affect the entire calculation consistently.The systematic evaluation of the predicted atomic data’ sensitivity to the optimised mean local potential provides valuable insights into the reliability and precision of the results.
@@ IV.1.3 Energy Levels
Figures 1 and 2 provide a comprehensive comparison of the energy levels for each angular momentum, , for Pt ii and Au ii, respectively. The benchmark data from NIST ASD are compared with the results of our calculations using FAC as well as the energy levels reported by Gillanders et al. using the GRASP0 code [18]. In particular, the comparison is done for each value and is not divided by parity, as the data obtained from Gillanders et al. did not include information on configurations or parity. This comparison highlights the effectiveness of the FMC optimisation in aligning our calculated energy levels with the NIST reference data, demonstrating significant improvements over the non-optimised results. For the non-optimised model, the electron occupancy number FMC was taken from an average using the ground and first configuration in energy, this is and for Pt ii and and for Au ii, applying the standard semi-empirical correction built into the code. In particular, for the lowest levels, the optimised results show a typical differences in excitation energy of less than 10% compared to the experimental data from NIST, in line with the results obtained for GRASP0 for the singly-ionised species. These results highlight the impact of the optimisation, particularly for low-lying levels with excitation energies below 7.5 eV, where large differences are found compared to the default FAC calculation. It is worth noting that the agreement between optimized calculations and NIST data for higher-lying states (above 7.5 eV) is better for Pt ii than for Au ii. This difference is partially attributable to the greater abundance of experimental data for Pt ii in the NIST ASD, as well as the more complex electronic structure of Au ii with its filled shell, which presents greater challenges for optimization due to stronger configuration mixing in higher-lying states.
Figure 3 illustrates the average relative difference to the data available in the NIST ASD for all identified levels of Pt ii and Au ii for different models computed with the FAC . The “No Opt.” model corresponds to an 8-configuration model that does not use the FAC potential optimisation, whereas the other models incorporate the optimisation. The impact of the optimisation is significant, resulting in a much greater improvement which was not achievable just by the inclusion of a larger CI basis. The results of the optimised models show a closer alignment to the NIST reference data, demonstrating similar accuracy to the GRASP0 [18] results and better energy than the HULLAC results [28, 55].
@@ IV.2.2 Computational procedure
The basis space for Pr ii and Pr iii was expanded to include all major configurations up to the ionisation energy. For Pr ii and Pr iii, it was found that considering single and double SD excitations from the ground configuration ( for Pr ii and for Pr iii) up to , , , , effectively covered the most relevant configurations that contribute directly to opacity [35]. These configurations were used in the optimisation procedure in order to determine the optimal FMC for these ions. Hence, we label this space as Optimisation Space (OPS). To further enhance the convergence of the results, additional configurations were included based on SD excitations extending up to principal quantum number ) angular momentum . The space was increased in multiple steps from a base ground configuration (GC), with each space used labelled as a different CI space (CIS). The scheme for the full set of calculations was as follows.
@@ IV.2.3 Energy Levels
The average relative difference from the NIST data for all identified levels of Pr ii and Pr iii for the different models calculated with FAC is shown in Fig. 6. The “No Opt.” model corresponds to the OPS configuration model that does not use the FAC potential optimisation, whereas the other models incorporate this optimisation. The optimisation leads to a substantial reduction in the relative difference, far exceeding the improvements achieved by CI alone. Additionally, the convergence of results across different CI spaces (CIS1, CIS2, and CIS3) indicates that the optimisation process stabilises the calculations, ensuring reliable and consistent results. The results for the energy levels of the first 20 computed levels are shown in Tables 1 and 2
@@ IV.2.4 Transitions
To evaluate the accuracy of our calculations for radiative transition probabilities and wavelengths in Pr ii and Pr iii, we compared our results with recent experimental data from Ferrara et al. (2024) [72]. Their study provides valuable spectroscopic data for several high- elements, including praseodymium, covering the 3700-10000 Å range using high-resolution échelle spectroscopy. Figure 7 presents a comparison between the calculated values from our CIS3 model and the experimental data. The experimental values are derived using two methods: Boltzmann plots and cascade calculations. It’s important to note that while cascade method results are generally reliable within the limits of adopted literature values, the Boltzmann method is only valid for transitions in thermodynamic equilibrium. When available, we also evaluated our data against data from the Vienna Atomic Line Database (VALD) [73]. As evident from Fig. 7, there is considerable scatter in the values, particularly for weaker transitions (). The agreement between our calculations and the experimental data improves for stronger transitions, but discrepancies persist across the entire range. Quantitatively, we observe a root mean square deviation (RMSD) of approximately 1.1 dex between our calculated values and the experimental data. This level of disagreement is not uncommon in atomic structure calculations for complex ions like Pr ii and Pr iii, especially given the challenges in accurately representing the electron correlations in these systems. In contrast, Fig. 8 shows good agreement between our calculated wavelengths and the experimental values. Most of the calculated wavelengths fall within 10% of the experimental data, as indicated by the points between the inner dotted lines. Nearly all calculated wavelengths are within 20% of the experimental values, falling between the dashed lines. Quantitatively, we find that approximately 80% of our calculated wavelengths deviate less than 20% from the experimental values and 96% are within 10%. This disparity in accuracy between the radiative transition probabilities and wavelengths is a common feature in atomic structure calculations. Wavelengths, which are primarily determined by energy level differences, are often more accurately predicted than transition probabilities, which depend on the precise details of the wavefunctions.
@@ IV.3.1 Available data
Building on this work, Goldschmidt (1963) [78] computed all 24 levels of the 6s configuration by fitting the Slater parameters using the initial eight experimentally determined levels as a reference. Subsequently, Marquet & Davis (1965) [79] identified 10 levels of the ground configuration using wavelengths in the visible and near-ultraviolet regions. Their Zeeman structure measurements of 132 transitions enabled Sluis & McNally (1970) [80] to identify 12 of the lowest ground configuration levels, which exhibited slightly higher energies than those calculated by Goldschmidt (1963) with = 30. Additionally, Sluis & McNally (1970) reported 90 odd levels, but these were excluded from comparisons due to a lack of configuration identification and assignment.
@@ IV.3.4 Transitions
In contrast, Fig. 13 demonstrates good agreement between our calculated wavelengths and the experimental values for Er ii. Most of the calculated wavelengths fall well within 10% of the experimental data, as indicated by the points that lie between the inner dotted lines. Virtually all calculated wavelengths are within 20% of the experimental values, falling between the dashed lines. Quantitatively, we find that approximately 85% of our calculated wavelengths deviate by less than 10% from the experimental values, and over 95% are within 20%. The good agreement in wavelength calculations for both Pr ii and Er ii ions represents a significant improvement over standard calculations without optimization and approaches the level of accuracy typically achieved with more computationally intensive multiconfiguration Dirac-Fock (MCDF) methods like those implemented in the GRASP code [30, 31]. This level of accuracy is particularly relevant for opacity calculations in kilonova modelling, while requiring significantly less computational resources than full MCDF approaches.
The stark contrast in accuracy between the radiative transition probabilities and the wavelengths observed for Er ii mirrors our findings for Pr ii. This discrepancy can be attributed to the nature of our optimisation procedure, which is primarily designed to minimise differences in energy levels when compared to experimental data. While this approach typically leads to more accurate wavefunctions, these results clearly demonstrate its limitations. The high accuracy in wavelength predictions, which are primarily determined by energy level differences, is a direct consequence of this optimisation strategy. However, persistent discrepancies in oscillator strengths suggest that our current approach may not fully capture the intricacies of the wavefunctions required for accurate transition probability calculations. This indicates that including a larger basis set in our calculations may be necessary to improve the accuracy of oscillator strengths, as these are more sensitive to the fine details of the electronic wavefunctions.
@@ VI Conclusions
However, persistent challenges in calculating accurate radiative transition probabilities underscore the need for further refinement of our approach. Future work should focus on expanding the basis set to potentially improve the accuracy of oscillator strengths and transition probabilities. Although computationally demanding, investigating the inclusion of core-core and core-valence correlations may further enhance the accuracy of energy level and transition calculations. While a perfect agreement for all atomic energy levels simultaneously is challenging due to the limitations of the local central potential approximation, our method provides a consistent improvement across a wide range of levels, and we also expect it produces more reliable wavefunctions supporting more accurate level identification. This enhanced identification capability provides a solid foundation for future work involving detailed calibration of theoretical energy levels against experimental data [89].